# Novel point cloud registration approach for noninvasive patient specific estimation of leaflet strain from 3D images of heart valves

### v1

@@ Abstract
Valvular heart disease is prevalent and a major contributor to heart failure. Valve leaflet strain is a promising metric for evaluating the mechanics underlying the initiation and progression of valvular pathology. However, robust and generalizable methods for noninvasively quantifying valvular strain from clinically acquired patient images remain limited. In this work, we present a novel feature-tracking framework for quantifying leaflet strain in atrioventricular valves using 3D echocardiographic images of pediatric and adult patients. Our method demonstrated superior accuracy in the assessment of anatomical deformation and strain of heart valves compared to other point-based approaches, as verified against a finite element benchmark. Further, our approach can robustly track inter-phase deformation of valves across highly variable morphologies without parameter tuning. Our analysis revealed that a median and interquartile range of the 1 principal strain greater than 0.5 is associated with leaflet billow (prolapse). Further investigation of the biomechanical signatures of heart valve disease has the potential to enhance prognostic assessment and longitudinal evaluation of valvular disease.
keywords:
@@
feature tracking, leaflet strain, cardiac valves, valvular heart disease1 Introduction
@@
Valvular heart disease (e.g., valve regurgitation or stenosis) is a prevalent cardiac condition that contributes substantially to the development of heart failure across all age groups [1, 2, 3, 4, 5]. Echocardiographic imaging is the clinical gold standard for evaluating morphological abnormalities of the heart valves (e.g., leaflet billowing, prolapse, and tenting) [6]. 3D echocardiography (3DE) in particular enables 3D visualization of dynamic valve geometry, as well as quantification of functional metrics such as regurgitant volume and effective orifice area, which are critical indicators of disease severity [7, 8, 9]. However, global morphological features alone are not sufficient to characterize the biomechanical environment that governs disease initiation, progression, or the long-term durability of surgical repair. Strain analysis, in contrast, provides a detailed map of local tissue deformation that reflects tissue stiffness. This information may serve as a proxy for identifying leaflet regions at increased risk of degeneration [10, 11, 12, 13]. Reliable non-invasive assessment of leaflet strain from 3D images thus holds great potential for advancing prognostic evaluation of valvular disease and improving predictions of surgical repair durability [14, 15].
Implications of leaflet strain on valve health and disease have been extensively investigated using finite element analysis (FEA) [16, 17, 18, 19, 20, 21]. While these seminal studies have provided important insights into the fundamental mechanisms of leaflet biomechanics, the fidelity of FEA is highly dependent on how accurately the model reproduces the native valvular system. To date, patient-specific FEA of heart valves remains limited by the spatial resolution of 3DE imaging and requires simplification and individualized adjustment of models, especially for chordal structures [16, 21]. Further, many of these studies also depend on leaflet tissue properties derived from animal or cadaver experiments, which further confounds simulation outcomes. The lack of a generalized framework for patient-specific valve modeling leads to user bias and variability, which may limit the reliability of these models in clinical practice.
@@
Feature tracking techniques, including image intensity–based registration [22, 23, 24] and point–based registration [25, 26], present an alternative approach for estimating leaflet strain. Feature tracking approximates local tissue deformation by aligning geometric features between image frames. This characteristic eliminates the need for prior knowledge of material properties or chordal structure, making this approach better suited for clinical translation. Notably, myocardial strain analysis has been widely adopted clinically to assess ventricular function. Multiple population studies suggested that myocardial strain analysis provides added value in improving diagnostic accuracy and prognostication in heart failure [27, 28, 29, 30, 31]. However, the application of registration-based approaches for valvular strain analysis remains in its infancy, with only a few notable studies to date [32, 15, 14]. This is primarily attributed to the challenge posed by large and rapid geometric changes in the leaflets during the transitions between open and closed time frames. While registration-based approaches hold considerable promise for valvular strain analysis, existing studies have focused exclusively on mitral valves (MVs), with most algorithms being proprietary or limited by reliance on commerical software designed for MVs alone [32, 15, 14].
@@
To address these limitations, we sought to develop a robust open-source, registration-based strain quantification framework for the assessment of diverse valve morphologies, including patients with congenital heart disease. Our method incorporates a novel distance-decay function into the Gaussian prior of the coherent point drift algorithm to robustly track geometries with substantial deformation [33, 26]. The proposed method was verified using an FEA benchmark example, where it achieved more accurate shape and strain estimation than other published point-based registration techniques. To further demonstrate robustness across a broad range of valve morphologies and types, we applied this verified framework to three patient cohorts, including adult and pediatric patients with congenital heart disease. Reliable feature tracking and strain quantification were achieved in both MVs and tricuspid valves (TVs) in patients across a broad age range, from infants to older adults with both normal and diseased morphologies.
@@ 2 Materials and Methods
Our proposed registration-informed strain quantification framework is presented in Fig. 1. In all subjects, the areal and 1 principal strain at the mid-systolic frame were computed using the mid-diastolic frame as reference. The code of our strain calculation framework will be made available at https://github.com/ww382/registration_informed_valve_strain.
2.1 Three-dimensional echocardiographic image processing
2.1.1 Atrioventricular valve reconstruction
@@
3DE images of pediatric TVs from patients with HLHS (n = 11; 2 TEE and 9 TTE) and MVs (n = 10; 3 TEE and 7 TTE) were obtained from an existing database at the Children’s Hospital of Philadelphia. These images were acquired on a Philips Epiq system (Philips Medical, Boston, MA). In addition, transesophageal 3DE images from adult patients (n = 10) with normal MV morphology were obtained from an existing database at the University of Pennsylvania and were acquired on a Philips iE33 system (Philips Medical Systems, Andover, MA). The pediatric data were imported into the SlicerHeart module [34, 9, 35] in 3D Slicer for segmentation, while the adult data were manually traced using ITK-SNAP [36]. The valve leaflets at the mid-diastolic and mid-systolic frames of the cardiac cycle were segmented by an expert observer to facilitate the 3DE-informed strain analysis. These were multi-label segmentations with a separate label for each leaflet. Image voxels corresponding to the leaflets were manually labeled and converted into 3D models (Fig. 1A), as described in [8, 7, 37, 38]. The collection and analysis of these data were approved by the Institutional Review Boards at the Children’s Hospital of Philadelphia and the University of Pennsylvania.
2.1.2 Atrial surface mesh creation
@@
The atrial surface of each leaflet segmentation was extracted using the Leaflet Analysis tool in SlicerHeart. In particular, a closed curve was defined along the margin of the leaflet that separates the atrial and ventricular sides of the segmentation model. The model was then clipped along this curve to isolate the atrial surface. A custom Python script that utilized the PyACVD library [39] was subsequently used to process the extracted surface. First, the extracted surface was cleaned by removing isolated and unreferenced points and then smoothed using Taubin smoothing (10 iterations, pass band = 0.05). Second, the surface was remeshed with 1,000 nodes uniformly distributed across the atrial surface.
2.2 Distance-weighted coherent point drift (CPD) registration
@@
The development of our DW-CPD registration approach was inspired by the original CPD formulation for nonrigid point set registration [26]. In the standard CPD framework, the fixed point set is represented as the centroids of a Gaussian mixture model (GMM) with isotropic covariances and uniform priors. The moving point set is transformed to the fixed point set through iterative expectation–maximization (EM) optimization to maximize the posterior correspondence probabilities (posterior probability function). The deformation field is regularized by the motion coherence theory [33], which penalizes non-smooth variations in the velocity field. This ensures that neighboring points maintain coherent movement and preserve the topological structure of the point set.
@@ 2.2.1 Gaussian mixture models
where is the initial nodal positions in the moving point set and is the displacement function.
@@ 2.2.2 Expectation-maximization optimization
CPD employs the EM algorithm to iteratively refine the nodal correspondences and registration parameters, as direct optimization of the log-likelihood is intractable. In the expectation (E) step, the posterior probability of each data point is given by
where is the matrix containing the probability of correspondence. In our DW-CPD approach, we modify this expression to prioritize point correspondences in neighboring points by introducing a distance decay function defined as
where is a distance decay factor, with a larger , the stronger the penalty is for enforcing the nodes in the moving frames to align very closely with the nodes in the fixed frame. Depending on the application, one may adjust to impose tighter (ensure node proximity) or looser (enables positional flexibility) locality constraints between the moving and fixed frame.
In the maximization (M) step, we solve for the matrix of the Gaussian kernel weights, , using the posterior probability function, , in the E step
where is a column vector of ones of length and is a symmetric Gram matrix with elements , where defines the width of the Gaussian filter. In the present work, we set = 1, = 5, and = 0.05.
@@
Here, we present our workflow for strain analysis on individual leaflets (Fig. 1B) as well as population-based analysis (Fig. 1C) based on nodal correspondences obtained from DW-CPD.
@@ 2.3.1 Individual lealfet analysis
Leaflet geometry in atrioventricular valves is highly heterogeneous, particularly in congenital heart disease [37]. To ensure robust registration, we first used the VTK Python module to align and rotate individual leaflets of the same label (e.g., anterior, posterior, septal) by their centroids so that their atrial surfaces face the same orientation. We manually checked the aligned leaflets to make sure the annular and free edges across all leaflets were in the same direction. A representative leaflet of diastole from one randomly selected patient was then chosen as the reference, and all other leaflets of the same anatomical category were subsequently registered to it using DW-CPD to establish intra-subject correspondences. From the registered leaflets, a mean shape was generated using generalized procrustes analysis. The mid-diastolic and mid-systolic point sets were then registered to this mean shape to establish point correspondences between the two phases. Finally, the deformation gradient and strain were computed from the nodal coordinates of the mid-diastolic and mid-systolic point sets.
Both areal strain and the 1 principal Green-Lagrange strain were considered in the present work. The areal strain of the mesh is formulated as
where and contain the area of each cell in its reference (mid-diastolic) and deformed (mid-systolic) state, respectively. The Green-Lagrange strain is formulated as
where is the right Cauchy-Green tensor and is the identify matrix. is formulated as where , the deformation gradient tensor, is computed as . Herein, and are the tangent vectors of each triangular cell in the reference state. The normal vector is computed as . and are the tangent vectors in the deformed state.
2.3.2 Population-based whole valve analysis
With one-to-one correspondence established across all leaflets, we applied generalized Procrustes analysis to iteratively align the valves and compute a mean shape for each patient cohort. Because of substantial variability in valve geometries, particularly within the HLHS TV cohort, some valves occasionally had misoriented leaflets, which could result in erroneous mean shape calculations. In such cases, the valves were manually rotated to the correct orientation. Principal component analysis was then performed to quantify shape variations at the mid-diastolic and mid-systolic frames. DW-CPD registration was performed on each pair of diastole-systole valve shapes, and strain was subsequently computed.

### v2

# A geometric feature tracking approach for noninvasive patient specific estimation of leaflet strain from 3D images of heart valves

@@ Abstract
Purpose Valvular heart disease is prevalent and a major contributor to heart failure. Valve leaflet strain is a promising metric for evaluating the mechanics underlying the initiation and progression of valvular pathology. However, generalizable methods for noninvasively quantifying valvular strain from clinically acquired patient images remain limited. This study aims to develop a robust feature-tracking framework that enables accurate shape matching across variable valve morphologies and quantification of in vivo atrioventricular leaflet strain from three-dimensional echocardiographic (3DE) images in pediatric and adult patients.
Methods We developed a geometric feature-tracking framework to quantify in vivo leaflet strain from 3DE images and to assess anatomical deformation across the cardiac cycle. The method integrates a cohort-derived geometric reference atlas to establish geometric correspondence and introduces a novel distance-weighted coherent point drift algorithm within a Gaussian mixture model framework for non-rigid registration. We evaluated performance against a finite element benchmark model and compared the approach with conventional point-based tracking methods. The framework was applied to pediatric and adult patient datasets (N = 31) to assess robustness across variable valve morphologies.
Results The proposed method demonstrated greater accuracy in quantifying anatomical alignment and leaflet strain than conventional point-based approaches. Validation against the finite element benchmark confirmed improved strain estimation. The framework achieved reliable inter-phase tracking of valve deformation across diverse morphologies in pediatric and adult patients. Analysis identified a consistent distribution pattern of the principal strain associated with leaflet billow (prolapse).
Conclusion This feature-tracking framework provides a generalizable method for noninvasive quantification of atrioventricular valve leaflet strain from clinical 3DE images. Characterization of biomechanical strain patterns may improve prognostic assessment and support longitudinal evaluation of valvular heart disease. Further investigation of the biomechanical signatures of heart valve disease has the potential to enhance prognostic assessment and longitudinal evaluation of valvular heart disease.
1 Department of Mechanical Engineering and Applied Mechanics, University of Pennsylvania, Philadelphia, PA.
2 Cardiovascular Institute, Children’s Hospital of Philadelphia, Philadelphia, PA.
3 Department of Anesthesiology and Critical Care Medicine, Children’s Hospital of Philadelphia, Philadelphia, PA.
4 Department of Biomedical Engineering, University of Utah, Salt Lake City, UT.
5 Scientific Computing Institute, University of Utah, Salt Lake City, UT.
6 Department of Radiology and Bioengineering, University of Pennsylvania, Philadelphia, PA.
@@
Keywords: cardiac valves, geometric feature tracking, leaflet strain analysis, valvular heart disease
1 Introduction
@@
Valvular heart disease (e.g., valve regurgitation or stenosis) is a prevalent cardiac condition that contributes substantially to the development of heart failure across all age groups [8, 6, 26, 36, 18]. Echocardiographic imaging is the clinical gold standard for evaluating morphological abnormalities of the heart valves (e.g., leaflet billowing, prolapse, and tenting) [24]. 3D echocardiography (3DE) in particular enables visualization of dynamic 3D valve geometry, as well as quantification of functional metrics such as regurgitant volume and effective orifice area, which are critical indicators of disease severity [30, 32, 35]. However, global morphological features alone are not sufficient to characterize the biomechanical environment that governs disease initiation, progression, or the long-term durability of surgical repair. Strain analysis, in contrast, provides a detailed map of local tissue deformation that reflects tissue stiffness and applied boundary conditions. This information may serve as a proxy for identifying leaflet regions at increased risk of degeneration and potential tissue damage or rupture [23, 55, 56, 42, 45]. Reliable non-invasive assessment of leaflet strain from 3DE images, transesophageal echocardiogram (TEE) and transthoracic echocardiogram (TTE), thus holds great potential for advancing prognostic evaluation of valvular disease and improving predictions of surgical repair durability [5, 10].
Assessment of leaflet strain is central to understanding the mechanical environment that drives tissue and cellular changes during valve disease initiation and progression [23, 55, 56, 42]. Finite element analysis (FEA) has been widely used to investigate leaflet biomechanics and to examine the effects of annular dynamics [25, 20], papillary muscle displacement [47], constitutive properties [52, 7], and annular geometry [43] on valve function [40, 34, 12, 25, 20, 22, 3, 37, 41]. These studies have provided important insights into the fundamental mechanisms of valvular biomechanics. Despite these advances, the fidelity of FEA depends on how accurately the model reproduces the native valvular system. Most studies rely on animal, cadaveric, or in vitro data, and many incorporate tissue properties derived from non-human specimens. Even with image-derived FEA that enables patient-specific reconstruction from clinical imaging [19, 40, 22, 3], significant challenges remain. The spatial resolution of 3DE limits geometric accuracy. Complex leaflet morphology complicates fitting of finite element structures to image-derived models. These approaches often require model simplification and individualized adjustment of models, particularly for subvalvular chordal structures. Assumptions regarding surrogate chordal force representations are frequently necessary. The absence of a standardized and generalized framework for patient-specific valve modeling introduces user bias and variability, which may limit clinical reliability.
@@
Feature tracking techniques, including image intensity–based feature tracking [13, 27, 16] and point–based feature tracking [4, 28], present an alternative approach for estimating leaflet strain. Feature tracking approximates local tissue deformation by aligning geometric features between image frames. This characteristic eliminates the need for prior knowledge of material properties or chordal structure, making this approach well-suited for clinical translation. Notably, myocardial strain analysis has been widely adopted clinically to assess ventricular function. Multiple population studies suggested that myocardial strain analysis provides added value in improving diagnostic accuracy and prognostication in heart failure [33, 11, 9, 57, 46]. While feature tracking approaches hold considerable promise for valvular strain analysis, application of feature tracking-based approaches for valvular strain analysis remains in its infancy, with only a few notable studies to date [2, 10, 5].
Existing valvular feature-tracking studies have focused exclusively on mitral valves (MVs), with most algorithms being proprietary or limited by reliance on commercial software designed for MVs alone [2, 10, 5]. Due to challenges posed by large and rapid geometric changes in the leaflets during the transitions between open and closed time frames, prior work primarily tracked valves from mid-systole to end-systole, which does not capture the near strain free diastolic state of an open valve as a baseline. While 3D + time (4D) CT [50] and cardiac magnetic resonance imaging [15] are increasingly being utilized to evaluate valve function, there is currently no generalizable framework for analysis of valve leaflet strain across all 3D modalities. The need for multimodality-based assessment of valve leaflet strain across diverse valve morphologies and populations, including patients with complex congenital heart disease, is the key motivation for the present study.
@@
To address these limitations, we sought to develop a robust open-source, feature tracking-based strain quantification framework for the assessment of diverse valve morphologies, including patients with congenital heart disease. Our method incorporates a novel distance-decay function into the Gaussian prior of the coherent point drift algorithm to robustly track geometries with substantial deformation [53, 28]. The proposed method was verified using an FEA benchmark example, where it achieved more accurate shape and strain estimation than other published point-based feature tracking techniques. To further demonstrate robustness across a broad range of valve morphologies and types, we applied this verified framework to three patient cohorts, including adult and pediatric patients with congenital heart disease. Reliable feature tracking and strain quantification were achieved in both MVs and tricuspid valves (TVs) in patients across a broad age range, from infants to older adults with both normal and diseased morphologies.
@@ 2 Materials and Methods
Our proposed feature tracking-informed strain quantification framework is presented in Fig. 1. In all subjects, the areal and 1 principal strain at the mid-systolic frame were computed using the mid-diastolic frame as reference. The code of our strain calculation framework will be made available at https://github.com/ww382/feature_tracking_informed_valve_strain.
2.1 Heart valve atrial surface processing
@@
Fig. 1A presents the process of preparing feature tracking-ready atrial surfaces of leaflets from 3DE images. 3DE images of pediatric TVs from patients with hypoplastic left heart syndrome (HLHS) (n = 11; 2 TEE and 9 TTE) and MVs (n = 10; 3 TEE and 7 TTE) were obtained from an existing database at the Children’s Hospital of Philadelphia. These images were acquired on a Philips Epiq system (Philips Medical, Boston, MA). In addition, transesophageal 3DE images from adult patients (n = 10) with normal MV morphology were obtained from an existing database at the University of Pennsylvania and were acquired on a Philips iE33 system (Philips Medical Systems, Andover, MA). Pre-existing segmentations for the authors’ research groups were utilized. Since the image data originated from labs at different institutions, which had separate protocols for image export and manual segmentation, the pediatric data were imported into the SlicerHeart module [44, 35, 21] in 3D Slicer for segmentation, while the adult data were manually traced using ITK-SNAP [54]. The valve leaflets at the mid-diastolic and mid-systolic frames of the cardiac cycle were segmented by an expert observer to facilitate the 3DE-informed strain analysis. These were multi-label segmentations with a separate label for each leaflet. Image voxels corresponding to the leaflets were manually labeled and converted into 3D models, as described in [32, 30, 31, 1]. The collection and analysis of these data were approved by the Institutional Review Boards at the Children’s Hospital of Philadelphia and the University of Pennsylvania.
@@
The atrial surface of each leaflet segmentation was extracted using the Leaflet Analysis tool in SlicerHeart. In particular, a closed curve was defined along the margin of the leaflet that separates the atrial and ventricular sides of the segmentation model. The model was then clipped along this curve to isolate the atrial surface. A custom Python script that utilized the PyACVD library [49] was subsequently used to process the extracted surface. First, the extracted surface was cleaned by removing isolated and unreferenced points and then smoothed using Taubin smoothing [48]. Second, the diastolic and systolic atrial surfaces of each leaflet were discretized to linear triangular meshes by distributing 1,000 nodes uniformly using Voronoi clustering.
2.2 Distance-weighted coherent point drift (DW-CPD) feature tracking
@@
The development of our DW-CPD feature tracking approach was inspired by the original CPD formulation for nonrigid point set feature tracking [28]. In the standard CPD framework, the fixed point set is represented as the centroids of a Gaussian mixture model (GMM) with isotropic covariances and uniform priors. The moving point set is transformed to the fixed point set through iterative expectation–maximization (EM) optimization to maximize the posterior correspondence probabilities (posterior probability function). The deformation field is regularized by the motion coherence theory [53], which penalizes non-smooth variations in the velocity field. This ensures that neighboring points maintain coherent movement and preserve the topological structure of the point set.
@@ 2.2.1 Gaussian mixture models
where is the initial nodal positions in the moving point set, is the deformation matrix of Gaussian kernel weights and is a symmetric Gram matrix with elements
and is the width of the Gaussian kernel and is the of the initial point cloud. The regularization function is expressed as
@@ 2.2.2 Expectation-maximization optimization
CPD formulates point set feature tracking as a probabilistic alignment problem by modeling one point set as centroids of a Gaussian mixture model and estimating a transformation that maximizes the likelihood of the observed data. Since direct optimization of the log-likelihood is intractable, CPD employs the expectation-maximization (EM) algorithm to estimate soft correspondences and feature tracking parameters iteratively.
In the expectation (E) step, given estimates of the transformation and the Gaussian variance , the posterior correspondence probability between the transformed source point and the target point is computed as
where denotes the soft posterior probability matrix and accounts for uniform noise and outliers.
In the proposed DW-CPD formulation, we introduce an additional distance-based weighting to encourage more localized correspondences. Specifically, the posterior probabilities are reweighted using an exponential distance decay function,
followed by normalization over for each . The decay parameter controls the strength of the locality bias, with larger values favoring short-distance correspondences and smaller values allowing greater positional flexibility.
In the maximization (M) step, the nonrigid transformation and the Gaussian variance are updated in a sequential manner. Given the previous estimate of the variance and the correspondence matrix , the deformation matrix of Gaussian kernel weights is updated by solving the following linear system
where is a column vector of ones of length , denotes the diagonal matrix whose diagonal entries are given by the vector , and is a regularization parameter enforcing motion coherence. This system admits a unique solution and yields an updated transformation defined in Equation 4.
With the updated transformation, the Gaussian variance is subsequently updated analytically as
which estimates the residual alignment error between the transformed source points and the target points.
The EM procedure alternates between estimating soft correspondences in the E step and sequentially updating the transformation and variance in the M step to iteratively refine the nonrigid feature tracking. In the present work, we set , , and . Furthermore, we assume no outliers in the point clouds by setting .
2.2.3 Feature tracking accuracy evaluation
Feature tracking accuracy was evaluated by quantifying the geometric agreement between the registered valve surfaces and the reference geometries. The mean symmetric distance (MSD) and the 95 %ile Hausdorff distance (HD95) were used to evaluate shape similarity. Let and denote the registered surface and the reference surface, respectively, each represented as a discrete set of surface points. For a point , the point-to-surface distance to is defined as
where denotes a point on the surface . The MSD is then computed as
HD95 is computed as,
where denotes the 95 %ile of the distribution.
@@ 2.2.3 Feature tracking accuracy evaluation
Here, we present our workflow for strain analysis on individual leaflets (Fig. 1B) as well as cohort-based analysis (Fig. 1C) based on nodal correspondences obtained from DW-CPD.
@@ 2.3.1 Individual lealfet analysis
Leaflet geometry in atrioventricular valves exhibits substantial inter-subject variability, particularly in congenital heart disease [31]. To enable consistent feature tracking across subjects and cardiac phases, we constructed a representative mean leaflet shape for each anatomical label. Individual leaflets of the same label (e.g., anterior, posterior, septal) were first aligned using the VTK Python module by translating and rotating each leaflet such that its centroid shared a common reference point and the atrial surfaces faced the same orientation. The aligned leaflets were manually inspected to ensure that annular and free edges were oriented consistently across samples. A representative mid-diastolic leaflet from a randomly selected subject was then chosen as a reference, and all other mid-diastolic leaflets of the same anatomical label were registered to this reference using DW-CPD to establish intra-subject point correspondences. From these registered shapes, a mean leaflet geometry was computed using generalized Procrustes analysis. The resulting mean shape serves as a reference atlas that facilitates interpolation and intermediate transitioning of leaflet geometry between the open and closed configurations. This intermediate representation is critical for robust feature tracking between mid-diastolic and mid-systolic frames, particularly in pathological valves (e.g., billowing leaflets), where large nonrigid deformations and pronounced changes in leaflet concavity occur. The mean shape is computed only once; thereafter, it is used to facilitate analysis of all patients’ valves.
The mid-diastolic and mid-systolic point sets were independently registered to the mean shape to establish consistent point correspondences across cardiac phases. The deformation gradient and strain were subsequently computed from the corresponding nodal coordinates between the two phases. Both areal strain and the 1 principal Green-Lagrange strain were considered in the present work. Areal strain provides an intuitive quantification of local in-plane surface expansion and compression during leaflet motion. In contrast, the 1 principal Green-Lagrange strain enables identification of regions experiencing elevated in-plane tensile deformation that may be susceptible to tissue damage or failure.
Areal strain. For a triangular element on the atrial surface with vertices , the surface is parameterized on a reference triangle using local coordinates and linear shape functions
which satisfy . The Cartesian coordinates of the surface are interpolated as
Here, denotes the vertex of a triangular element, written as a column vector .
The surface Jacobian is defined as
where the derivatives of the shape functions are constant over each element. The surface Jacobian yield two surface tangent vectors,
The element area is then evaluated as
For each triangulated surface element, the surface area change is quantified using the areal strain,
where and denote the element areas in the reference (mid-diastolic) and deformed (mid-systolic) configurations, respectively.
Green-Lagrange strain. To compute in-plane strain, a local deformation gradient is estimated by mapping the reference element basis to the corresponding deformed element basis. For each triangular element, two surface tangent vectors, and , are first obtained from the surface Jacobian. The element normal is then computed as
Using the tangent and normal vectors, the surface basis matrices for the reference and deformed configurations are constructed as
where and are the surface tangents obtained from the Jacobian, and is the corresponding unit normal.
The deformation gradient mapping the reference configuration to the deformed configuration is computed as
The right Cauchy–Green tensor and Green-Lagrange strain tensor are then defined as
where is the identity tensor.
2.3.2 Cohort-based whole valve analysis
With point-to-point correspondence established across all valves, we performed a cohort-based shape analysis to characterize morphological variability within each patient cohort and to provide geometric context for the interpretation of leaflet strain. The objectives of this analysis were to construct statistical shape models of atrioventricular valve geometry at mid-diastole and mid-systole, quantify the dominant modes of geometric variation within each cohort, and relate these variations to observed strain patterns.
Valve geometries at each cardiac phase (mid-diastole and mid-systole) were analyzed separately. Within each phase, all valve shapes were rigidly aligned using generalized Procrustes analysis such that they share the same origin, orientation, and scale. A mean valve geometry was then computed for each cohort and cardiac phase. Owing to substantial inter-subject variability in valve anatomy and orientation, particularly within the HLHS tricuspid valve cohort, automatic alignment occasionally resulted in inconsistent leaflet orientations. In such cases, the misaligned valves were manually rotated to ensure consistent anatomical orientation prior to mean shape computation.
Principal component analysis (PCA) was subsequently performed on the aligned valve geometries at each cardiac phase to quantify shape variability within each patient cohort. Each valve shape was represented as a high-dimensional vector formed by concatenating the three-dimensional coordinates of all mesh nodes. The resulting principal components define orthogonal modes of geometric variation about the mean shape. Shape variation along each mode was achieved by reconstructing shapes at and standard deviations (SD) from the mean according to
Here, denotes the number of standard deviations along the principal component used to reconstruct shape variation, denotes the mean shape, and and are the eigenvalue and eigenvector associated with the principal component.
To evaluate leaflet deformation and strain between cardiac phases, diastolic and systolic valve geometries for each subject were then registered using DW-CPD, and areal and Green-Lagrange strains were computed on the resulting corresponding meshes. As a result, strain patterns can be further analyzed within the context of inter-subject morphological variability.